OpenStudy (anonymous):

(ANSWER PLEASE)Let sin x = .72 and cos y = .72. Which is csc x – csc y? A. .65 B. -.56 C. .60 D. -.06

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

What is it? @Tehsh

OpenStudy (anonymous):

no the equation is correctly written as above.

OpenStudy (anonymous):

Here it goes, first we have to find x so we derive that out of sin x=.72. First step is to get x alone. To do that we have to multiply both side by the inverse of sin also known as \[\sin^{-1} \]. When you multiply both sides by sin^-1 you get \[x=\sin^{-1} (.72)\] Plug that into your calculator and you get \[.8038\]. If you don't know how to plug in \[\sin^{-1} \] First press the 2nd button on your calculator and then sin and it gives you the inverse. Remember to be in radians mode when you plug in the inverse of sin. So the next step is to find y. Just like x we have to get y by itself so we divide by the inverse of cos and we get\[y=\cos^{-1} (.72)\] When you plug that into the calc you get \[.7669\] Now you have to remember the trig rules. Remember that csc is the inverse of of sin and so cscx-cscy is the same thing as \[\frac{ 1 }{ \sin x }-\frac{ 1 }{ \sin y }\] So since we have x and y now plug that into the equation\[\frac{ 1 }{\sin( .8038) }-\frac{ 1 }{ \sin (.7669) }\] simplify that out and you will get that the answer

OpenStudy (anonymous):

No the first equation was rightt

OpenStudy (anonymous):

\[\sin^2y= 1-\cos^2y=1-(0.72)^2= 1-0.5184=0.4816\] \[siny= \sqrt{0.4816}=0.6939\approx0.69\] \[cosec y= \frac{1}{siny}=\frac{1}{0.69}\] \[cosec x= \frac{1}{sinx}=\frac{1}{0.72}\] \[cosecx- cosecy=\frac{1}{0.72}-\frac{1}{0.69}\] \[\[cosecx- cosecy=\frac{0.69-0.72}{0.72\times0.69}=\frac{-0.03}{0.4968}=-0.06038\approx-0.06\]

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

So option D. -.06 is correct

OpenStudy (anonymous):

Excelente!

OpenStudy (anonymous):

Thank you guys ....